26) Suppose the L and M genes are on the same chromosome but separated by 100 map units. What fraction of the progeny from the cross L M / l m × l m / l m would be L m / l m?
Answer #1

recombination frequency in percentage is the distance between genes in map units.

here distance between genes in map units is 100 map units,

the maximum recombination frequency is 50%, if the recombination frequency between the genes is more than 50% the genes are not linked on a chromosome.

here the distance between the genes is 100 map units so the genes are not linked.

LM / lm * lm/lm

LM Lm lM lm
lm LM/lm Lm/lm lM/lm lm/lm

the expected ratio is

LM/lm:Lm/lm:lM/lm:lm/lm=1:1:1:1

all the genotypes are expected in the equal proportion=1/4 that is 25% each

fraction of the progeny from the cross L M / l m × l m / l m would be L m / l m= 1/4